# Adapting a novel framework for rapid inference of massive black hole binaries for LISA

### v1

@@ Abstract
The Laser Interferometer Space Antenna (LISA) is designed to detect a variety of gravitational-wave events, including mergers of massive black hole binaries, stellar-mass black hole inspirals, and extreme mass-ratio inspirals. LISA’s capability to observe signals with high signal-to-noise ratios raises concerns about waveform accuracy. Additionally, its ability to observe long-duration signals will raise the computational cost of Bayesian inference, making it challenging to use costly and novel models with standard stochastic sampling methods without incorporating likelihood and waveform acceleration techniques. In this work, we present our attempt to tackle these issues. We adapt RIFT for LISA to take advantage of its embarrassingly parallel architecture, enabling efficient analysis of large datasets with costly gravitational wave models without relying on likelihood or waveform acceleration. We demonstrate that we can accurately infer parameters of massive black hole binary signals by carrying out a zero-noise injection recovery using the numerical relativity surrogate model NRHybSur3dq8. By utilizing all available $m\neq 0$ modes in the inference, we study the impact of higher modes on LISA data analysis. We study the impact of multiple massive black hole binary signals in a dataset on the inference of a single signal, showing that the selected source’s inference remains largely unaffected. Furthermore, we analyze the LDC-1A and blind LDC-2A datasets from the Radler and Sangria challenge of the LISA data challenges. When eschewing specialized hardware, we find NRHybSur3dq8 injection-recovery takes approximately $20$ hours to complete, while the analysis of Sangria and Radler datasets takes about $10$ hours to complete.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017) presents exciting scientific opportunities by operating in the millihertz frequency range, enabling the detection of sources currently inaccessible to ground-based observatories like the LIGO-Virgo-KAGRA network The LIGO Scientific Collaboration et al. (2015); Acernese et al. (2015); Akutsu et al. (2021). Among these sources are massive black hole binaries (MBHBs) Klein et al. (2016), stellar-mass black hole inspirals Sesana (2016), extreme mass ratio inspirals (EMRIs) Glampedakis et al. (2002); Babak et al. (2007, 2017), and galactic white dwarf binaries (GBs) Nelemans, G. et al. (2001). While the detection of this diverse array of signals promises exciting scientific insights, analyzing data from LISA poses challenges not faced by ground-based detectors.
@@
LISA is anticipated to detect a large number of signals Amaro-Seoane et al. (2023), with durations ranging from days to years. The combination of long signal durations and a high detection rate will lead to overlapping signals, necessitating simultaneous characterization of gravitating sources through global fitting pipelines Littenberg and Cornish (2023); Katz et al. (2024); Strub et al. (2024). Moreover, such large datasets will require efficient analysis algorithms to manage the increased computational cost of Bayesian inference. Furthermore, some of these signals will possess significantly high signal-to-noise ratios (SNR), which will require accurate waveform models Group et al. (2023) to be used in inference to avoid biases that could influence scientific conclusions. Other challenges include accurately modeling the LISA response for unbiased inference of source parameters. Moreover, the data is also expected to contain non-stationary and non-Gaussian noise, as well as glitches and gaps, all of which can complicate analysis. Successfully addressing these challenges is essential for maximizing LISA’s scientific output.
@@
The primary focus of the LISA mission is the detection of MBHBs, with recent studies indicating a detection rate of $1$ to $20$ per year Klein et al. (2016); Berti et al. (2016); Salcido et al. (2016); Katz et al. (2019); Bonetti et al. (2019). MBHBs are anticipated to offer valuable insights into a range of astrophysical, cosmological, and fundamental physics questions Arun et al. (2022), with SNRs ranging from hundreds to thousands. Although such high SNR signals will be rich in scientific information, the accuracy of the extracted information will be limited by the finite accuracy of waveform models and the numerical relativity waveforms used in their construction.
@@
Given the importance of MBHBs, considerable effort has been dedicated to analyzing them with prior studies utilizing the Fisher matrix approach Cutler (1998); Vecchio (2004); Berti et al. (2005); Lang and Hughes (2007); Arun et al. (2009) to estimate statistical uncertainties in MBHB parameters. However, this approach by construction cannot handle multimodal distributions, is only applicable in high SNR limit, and has been shown to not encompass necessary information to make general statements about parameter estimation of MBHBs Porter and Cornish (2015, 2008); Vallisneri (2008). As such the community has made a collective effort to develop and utilize Bayesian inference tools. The first study to carry out fully Bayesian analyses on inspiral, merger, ringdown MBHB signals (albeit non-spinning MBHBs) was Marsat et al. (2021). Since then numerous works have analyzed complete inspiral, merger, ringdown MBHB signals using Bayesian methods. In Cornish and Shuman (2020); Katz (2022) a complete analysis of the LISA Data Challenge (LDC) dataset, LDC-1A, including the initial search was carried out. The work in Katz et al. (2020) expanded upon the work of Marsat et al. (2021) to include aligned spins. Different groups presented their solutions to the LDC-2A Sangria dataset by using global fit Littenberg and Cornish (2023); Katz et al. (2024); Strub et al. (2024), template bank search followed by inference Weaving et al. (2023) and inference of a single signal Hoy and Nuttall (2023). All these Bayesian analyses used phenomenological models Ajith et al. (2007, 2008, 2011) to generate template waveforms, except Marsat et al. (2021) which used a frequency domain Reduced-Order-Model of EOBNRv2HM Pan et al. (2011). Also, most of these studies benefited from likelihood acceleration techniques like heterodyning Cornish (2010); Zackay et al. (2018); Cornish (2021). While heterodyning can speed up analysis by a factor of $10^{2}$ Cornish (2021), it requires knowledge of a high likelihood fiducial point and can produce an observable bias in the recovered posterior distributions.
@@
While theoretical studies have explored the effects of inaccuracies in waveform models Cutler and Vallisneri (2007) and numerical relativity waveforms Ferguson et al. (2021) on MBHB data analysis, there has been limited work assessing this impact using a full Bayesian approach. Therefore, in this paper, we adapt RIFT Pankow et al. (2015); Lange et al. (2018, 2017); Wysocki et al. (2019); Wofford et al. (2023) a highly parallelizable, grid-based parameter estimation code for LISA, henceforth called LISA-RIFT Jan (2024). RIFT is one of the several parameter inference algorithms that have been effectively utilized to infer binary system parameters from gravitational wave (GW) observations The LIGO Scientific Collaboration et al. (2019, 2021a, 2021b, 2021c). Its main advantage lies in its ability to rapidly and accurately infer parameters using costly or novel models, without relying on techniques to accelerate waveform generation Pürrer (2014) or likelihood evaluation Morisaki (2021); Cornish (2010); Antil et al. (2013); Canizares et al. (2013, 2015); Zackay et al. (2018); Leslie et al. (2021); Cornish (2021), which often serves as a bottleneck for stochastic sampling methods. RIFT’s unique approach also facilitates the study of waveform systematics by enabling direct likelihood comparisons, and model selection by generating low-cost, high-accuracy model evidence, making it an ideal tool for extensive waveform systematics studies for LISA.
@@
In this paper, we discuss our modifications to RIFT to carry out inference of LISA signals, including the implementation of the full time-frequency dependent LISA instrument response. We show that LISA-RIFT can accurately infer MBHB parameters by recovering a multi-mode signal generated using NRHybSur3dq8 Varma et al. (2019) at zero-noise realization. Additionally, we investigate how the presence of multiple MBHB signals in a dataset affects the inference of a single signal. We also analyze the LDC-1A and LDC-2A datasets from the Radler and Sangria challenge Babak and Petiteau ; Baghi (2022), discussing the impact of instrumental and galactic binary foreground noise on MBHB inference.
@@
The paper is organized as follows: in Sec. II, we briefly outline RIFT’s approach to parameter estimation, describe the implemented LISA instrument response, detail improvements in the initial grid generation, and discuss efforts to enhance user-friendliness. In Sec. III, we present the results from analyzing a multi-mode zero-noise MBHB signal using NRHybSur3dq8, our study on the impact of multiple MBHB signals on the inference of a single signal, and the analysis of zero-noise LDC-1A and noisy LDC-2A datasets. In Sec. IV, we discuss LISA-RIFT’s performance. Finally, in Sec. V we summarize our main findings and suggest directions for future research.
@@ II.1 Brief overview
Here, $p(\bm{\theta})$ represents the priors over $\bm{\theta}$ variables. This integration is made possible by expressing GW signals in terms of modes $h_{\ell m}(m_{1},m_{2},\bm{a_{1}},\bm{a_{2}};t)$, which are associated with the spin-weighted spherical harmonic decomposition of radiation in all possible directions. The plus $h_{+}(t)$ and the cross polarizations of the GW $h_{\times}(t)$ are related to the $h_{\ell m}(t)$ as:
@@ II.2 LISA response
Here, $\Psi_{\ell m}(f)$ is the phase of the frequency domain mode $h_{\ell m}(f)$. These single-link observables are affected by laser noise, which is several orders of magnitude larger than the signal. To address this, time-delay interferometry (TDI) is employed, allowing the use of these observables to construct a new set of virtual interferometric observables with significantly suppressed laser frequency noise. The LDC datasets provide three such TDI observables: $X,Y,Z$ Tinto and Armstrong (1999); Armstrong et al. (1999); Estabrook et al. (2000); Tinto (2000); Dhurandhar et al. (2002). However, these observables exhibit correlated noise properties and therefore, in this work, we use the uncorrelated combinations $A,E,$ and $T$. These observables are related to the TDI observables $X,Y,Z$ as follows Vallisneri (2005):
@@
We validated our implementation of the response against BBHx Katz et al. (2020); Katz (2022, 2021). We found our implementation has an excellent agreement with BBHx, with mismatches less than $10^{-10}$. For generating GW modes, we utilize the existing RIFT waveform interface, allowing the use of all implemented waveform models for data analysis. In this work, we used the LALSimulation library to generate modes for NRHybSur3dq8 Varma et al. (2019) and IMRPhenomD Khan et al. (2016).
@@ II.3 Well placed initial grid
The first step in a run is to set up an initial grid, which requires prior knowledge of a point with significant likelihood and the SNR. Ideally, this information comes from a search pipeline. Using this information, we create a grid on chirp mass ${\cal M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$, symmetric mass ratio $\eta=(m_{1}m_{2})/(m_{1}+m_{2})^{2},~{}a_{1z},~{}a_{2z},~{}\beta$ and $\lambda$. We estimate the bounds of this grid by calculating the Fisher errors using the 2-PN equation from Poisson and Will (1995). We also multiply the estimates for each parameter by a factor ($>1$) to address possible systematics and to compensate for the tendency of these equations to underestimate statistical errors for any given signal. We found that constructing a regular grid for ${\cal M}_{c}$ and $\eta$, while randomly sampling points from the ranges of $a_{1z},a_{2z},\beta$ and $\lambda$ leads to the fastest convergence. The multiplicative factor on each parameter ensures that the grid is compact enough to avoid unnecessary exploration in low-likelihood regions, yet broad enough to enclose the entire posterior distribution and account for potential systematics.
@@ III Results
We employ the AV sampler (based on the Varaha sampler Tiwari et al. (2023); Tiwari (2024)) to carry out adaptive importance sampling to evaluate marginalized likelihood and to construct posteriors. Since likelihood evaluation is not a bottleneck, we calculate the full Whittle likelihood artist Moran and Whittle (1951) over the frequency range of $0.1$ mHz to $0.1$ Hz at the full Fourier bin width. However, for our analysis of the LDC-1A dataset we integrate up to $0.05$ Hz, which is the Nyquist frequency of that dataset. We confirm the convergence of the algorithm by ensuring that the Jensen-Shannon divergence Lin (1991) between the posteriors from the last two iterations is below $0.005$. All the posteriors presented here were generated using $20,000$ samples.
@@ III.1 Zero-noise injection
We inject an MBHB signal, with parameters summarized in Table 1, into approximately four months of data sampled at five-second intervals at a zero-noise realization. The signal was generated using numerical relativity surrogate model NRHybSur3dq8 allowing us to incorporate a much larger set of higher-order modes in our analysis compared to previous studies Marsat et al. (2021); Katz et al. (2020), with the injected signal including content from the following modes: (2,2),(3,3),(4,4),(5,5),(2,1),(3,2),(3,1),(4,3) and (4,2). The inference was performed using the same model and with the same mode content. This was done to eliminate the impact of model waveform systematics and isolate the impact of higher modes on data analysis. Both the injected waveform and recovery templates were generated at a $f_{\text{min}}$ of $0.6$ mHz. To ensure higher-order modes are excited and significantly impact SNR, we set the mass ratio of the injection to approximately $1/3$, observed at an $\iota$ of $\pi/4$. The distance was chosen such that the signal will be observed at a high SNR of 1000. For likelihood computations, we use the lisa-data-challenge working group python package to generate the power spectral density (PSD), with the specific noise model used being “SciRDv1”.
@@
Figure 2: Zero-noise injection recovery using NRHybSur3dq8: This figure illustrates the results of our analysis of the zero-noise NRHybSur3dq8 injection. For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. In the top plot, the colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values. For clarity, the reflected sky mode is not included in this plot.
@@
The results of our analysis of this mock signal are presented in Fig 2. Due to the high SNR and the presence of higher mode content, the posteriors for mass and spins are Gaussian-like and peak at the injected values. The typical bias observed in sky localization for (2,2) only mode analyses Marsat et al. (2021) is absent, with the posteriors peaking at the true values. Furthermore, when considering the full time and frequency dependence of the LISA response and conducting a complete inspiral-merger-ringdown analysis, only the true and reflected modes [^1] The reflected sky mode (*) is related to the true sky mode through the following transformations in the LISA ($L$) frame:          $\beta^{*}_{L}=-\beta_{L}$, $\lambda^{*}_{L}=\lambda_{L}$. for sky position are observed. The reflected mode for this signal, located at $\beta=0.436$ and $\lambda=1.199$, has a low probability; for clarity, we exclude the reflected sky mode from Fig 2. Additionally, we note that the degeneracy between $D_{L}$ and $\iota$ is broken, resulting in two Gaussian-like posteriors that peak at the injected values. In contrast, in Sec. III.2 where the injection and recovery is performed using a (2,2) only mode model, we begin to observe biases in sky location, $D_{L}$, and $\iota$.

### v2

@@ Abstract
The Laser Interferometer Space Antenna (LISA) is designed to detect a variety of gravitational-wave events, including mergers of massive black hole binaries, stellar-mass black hole inspirals, and extreme mass-ratio inspirals. LISA’s capability to observe signals with high signal-to-noise ratios raises concerns about waveform accuracy. Additionally, its ability to observe long-duration signals will raise the computational cost of Bayesian inference, making it challenging to use costly and novel models with standard stochastic sampling methods without incorporating likelihood and waveform acceleration techniques. In this work, we present our attempt to tackle these issues. We adapt RIFT for LISA to take advantage of its embarrassingly parallel architecture, enabling efficient analysis of large datasets with costly gravitational wave models without relying on likelihood or waveform acceleration. We demonstrate that our open-source code can accurately infer parameters of massive black hole binary signals by carrying out a zero-noise injection recovery using the numerical relativity surrogate model NRHybSur3dq8. By utilizing all available $m\neq 0$ modes in the inference, we study the impact of higher modes on LISA data analysis. We study the impact of multiple massive black hole binary signals in a dataset on the inference of a single signal, showing that the selected source’s inference remains largely unaffected. Furthermore, we analyze the LDC-1A and blind LDC-2A datasets from the Radler and Sangria challenge of the LISA data challenges. When eschewing specialized hardware, we find NRHybSur3dq8 injection-recovery takes approximately $20$ hours to complete, while the analysis of Sangria and Radler datasets takes about $10$ hours to complete.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017) presents exciting scientific opportunities by operating in the millihertz frequency range, enabling the detection of sources currently inaccessible to ground-based observatories like the LIGO-Virgo-KAGRA network The LIGO Scientific Collaboration et al. (2015); Acernese et al. (2015); Akutsu et al. (2021). Among these sources are massive black hole binaries (MBHBs) Klein et al. (2016), stellar-mass black hole inspirals Sesana (2016), extreme mass ratio inspirals (EMRIs) Glampedakis et al. (2002); Babak et al. (2007, 2017), and galactic white dwarf binaries (GBs) Nelemans, G. et al. (2001). Ground-based detectors are unable to detect these low-frequency signals due to limitations imposed by factors such as seismic noise Martynov et al. (2016) and the restricted size of the detectors, challenges that LISA is designed to overcome.
@@
While LISA’s ability to detect a diverse array of signals promises exciting scientific insights, it also introduces challenges not encountered with ground-based detectors. LISA is anticipated to detect a large number of signals Amaro-Seoane et al. (2023), with durations ranging from days to years. The combination of long signal durations and a high detection rate will lead to overlapping signals, necessitating simultaneous characterization of gravitating sources through global fitting pipelines Littenberg and Cornish (2023); Katz et al. (2024); Strub et al. (2024). Moreover, such large datasets will require efficient analysis algorithms to manage the increased computational cost of Bayesian inference. Furthermore, some of these signals will possess significantly high signal-to-noise ratios (SNR), which will require accurate waveform models Group et al. (2023) to be used in inference to avoid biases that could influence scientific conclusions. Other challenges include accurately modeling the LISA response for unbiased inference of source parameters. Moreover, the data is also expected to contain nonstationary and non-Gaussian noise, as well as glitches and gaps, all of which can complicate analysis. Successfully addressing these challenges is essential for maximizing LISA’s scientific output.
@@
The primary focus of the LISA mission is the detection of MBHBs, with recent studies indicating a detection rate of $1$ to $20$ per year Klein et al. (2016); Berti et al. (2016); Salcido et al. (2016); Katz et al. (2019); Bonetti et al. (2019). MBHBs are anticipated to offer valuable insights into a range of astrophysical, cosmological, and fundamental physics questionsi Arun et al. (2022), with SNRs ranging from hundreds to thousands. Although such high SNR signals will be rich in scientific information, the accuracy of the extracted information will be limited by the finite accuracy of waveform models and the numerical relativity waveforms used in their construction.
@@
Given the importance of MBHBs, considerable effort has been dedicated to analyzing them with prior studies utilizing the Fisher matrix approach Cutler (1998); Vecchio (2004); Berti et al. (2005); Lang and Hughes (2007); Arun et al. (2009) to estimate statistical uncertainties in MBHB parameters. However, this approach by construction cannot handle multimodal distributions, is only applicable in high SNR limit, and has been shown to not encompass necessary information to make general statements about parameter estimation of MBHBs Porter and Cornish (2015, 2008); Vallisneri (2008). As such the community has made a collective effort to develop and utilize Bayesian inference tools. The first study to carry out fully Bayesian analyses on inspiral, merger, ringdown MBHB signals (albeit nonspinning MBHBs) was Marsat et al. (2021). Since then numerous works have analyzed complete inspiral, merger, ringdown MBHB signals using Bayesian methods. In Cornish and Shuman (2020); Katz (2022) a complete analysis of the LISA Data Challenge (LDC) dataset, LDC-1A, including the initial search was carried out. The work in Katz et al. (2020) expanded upon the work of Marsat et al. (2021) to include aligned spins. Different groups presented their solutions to the LDC-2A Sangria dataset by using global fit Littenberg and Cornish (2023); Katz et al. (2024); Strub et al. (2024), template bank search followed by inference Weaving et al. (2023) and inference of a single signal Hoy and Nuttall (2023). The analysis in Hoy et al. (2024) significantly decreased inference time by making intuitive arguments to constrain the search parameter space consistent with the observed data. All these Bayesian analyses used phenomenological models Ajith et al. (2007, 2008, 2011) to generate template waveforms, except Marsat et al. (2021) which used a frequency domain reduced-order-model of EOBNRv2HM Pan et al. (2011). Also, most of these studies benefited from likelihood acceleration techniques like heterodyning Cornish (2010); Zackay et al. (2018); Cornish (2021). While heterodyning can speed up analysis by a factor of $10^{2}$ Cornish (2021); Hoy and Nuttall (2023), it requires knowledge of a high likelihood fiducial point and can produce an observable bias in the recovered posterior distributions, even when the true value is used as the fiducial point Hoy and Nuttall (2023).
@@
While theoretical studies have explored the effects of inaccuracies in waveform models Cutler and Vallisneri (2007) and numerical relativity waveforms Ferguson et al. (2021) on MBHB data analysis, there has been limited work assessing this impact using a full Bayesian approach. Therefore, in this paper, we adapt RIFT Pankow et al. (2015); Lange et al. (2018, 2017); Wysocki et al. (2019); Wofford et al. (2023) a highly parallelizable, grid-based parameter estimation code for LISA, henceforth called LISA-RIFT (open-source code repository Jan (2024)). RIFT is one of the several parameter inference algorithms that have been effectively utilized to infer binary system parameters from gravitational wave (GW) observations The LIGO Scientific Collaboration et al. (2019, 2021a, 2021b, 2021c). Its main advantage lies in its ability to rapidly and accurately infer parameters using costly or novel models, without relying on techniques to accelerate waveform generation Pürrer (2014) or likelihood evaluation Morisaki (2021); Cornish (2010); Antil et al. (2013); Canizares et al. (2013, 2015); Zackay et al. (2018); Leslie et al. (2021); Cornish (2021), which often serves as a bottleneck for stochastic sampling methods. RIFT’s unique approach also facilitates the study of waveform systematics by enabling direct likelihood comparisons, and model selection by generating low-cost, high-accuracy model evidence, making it an ideal tool for extensive waveform systematics studies for LISA.
@@
In this paper, we discuss our modifications to RIFT to carry out inference of LISA signals, including the implementation of the full time-frequency dependent LISA instrument response. We show that LISA-RIFT can accurately infer MBHB parameters by recovering a multimode signal generated using NRHybSur3dq8 Varma et al. (2019) at zero-noise realization. Additionally, we investigate how the presence of multiple MBHB signals in a dataset affects the inference of a single signal. We also analyze the LDC-1A and LDC-2A datasets from the Radler and Sangria challenge Babak and Petiteau ; Baghi (2022), discussing the impact of instrumental and galactic binary foreground noise on MBHB inference.
@@
The paper is organized as follows: in Sec. II, we briefly outline RIFT’s approach to parameter estimation, describe the implemented LISA instrument response, detail improvements in the initial grid generation, and discuss efforts to enhance user-friendliness. In Sec. III, we present the results from analyzing a multimode zero-noise MBHB signal using NRHybSur3dq8, our study on the impact of multiple MBHB signals on the inference of a single signal, and the analysis of zero-noise LDC-1A and noisy LDC-2A datasets. In Sec. IV, we discuss LISA-RIFT’s performance. Finally, in Sec. V we summarize our main findings and suggest directions for future research.
@@ II.1 Brief overview
Here, $p(\bm{\theta})$ represents the priors over $\bm{\theta}$ variables. This integration is made possible by expressing GW signals in terms of modes $h_{\ell m}(m_{1},m_{2},\bm{a_{1}},\bm{a_{2}};t)$, which are associated with the spin-weighted spherical harmonic decomposition of radiation in all possible directions. The plus $h_{+}(t)$ and the cross polarizations $h_{\times}(t)$ of the GW are related to $h_{\ell m}(t)$ as:
@@ II.2 LISA response
Here, $\Psi_{\ell m}(f)$ is the phase of the frequency domain mode $h_{\ell m}(f)$. These single-link observables are affected by laser noise, which is several orders of magnitude larger than the signal Tinto and Dhurandhar (2005). To address this, time delay interferometry (TDI) is employed, allowing the use of these observables to construct a new set of virtual interferometric observables with significantly suppressed laser frequency noise. The LDC datasets provide three such TDI observables: $X,Y,Z$ Tinto and Armstrong (1999); Armstrong et al. (1999); Estabrook et al. (2000); Tinto (2000); Dhurandhar et al. (2002). However, these observables exhibit correlated noise properties and therefore, in this work, we use the uncorrelated combinations $A,E,$ and $T$. These observables are related to the TDI observables $X,Y,Z$ as follows Vallisneri (2005):
@@
We validated our implementation of the response against BBHx Katz et al. (2020); Katz (2022, 2021). We found our implementation has an excellent agreement with BBHx. To quantify this, we computed mismatches between our response and response obtained using BBHx and found the mismatches to be of $O(10^{-10})$. Mismatches range between 00 and $1$, where a mismatch of 00 indicates that the two series are identical. At this negligible level of mismatch, any bias introduced would only become significant Lindblom et al. (2008) at an SNR significantly higher than what LISA is expected to observe. For generating GW modes, we utilize the existing RIFT waveform interface, allowing the use of all implemented waveform models for data analysis. In this work, we used the LALSimulation LIGO Scientific Collaboration et al. (2018) library to generate modes for NRHybSur3dq8 Varma et al. (2019) and IMRPhenomD Khan et al. (2016).
@@ II.3 Well placed initial grid
The first step in a run is to set up an initial grid, which requires prior knowledge of a point with significant likelihood and the SNR. The choice of this point does not affect the final posterior distributions. Ideally, this information is obtained from a search pipeline, which, for LISA, could also be derived from a global fit pipeline. Using this information, we create a grid on chirp mass ${\cal M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$, symmetric mass ratio $\eta=(m_{1}m_{2})/(m_{1}+m_{2})^{2},~{}a_{1z},~{}a_{2z},~{}\beta$ and $\lambda$. We estimate the bounds of this grid by calculating the Fisher errors using the 2-PN equation from Poisson and Will (1995). We also multiply the estimates for each parameter by a factor ($>1$) to address possible systematics and to compensate for the tendency of these equations to underestimate statistical errors for any given signal. For each of our runs, we construct a regular grid on ${\cal M}_{c}$ and $\eta$, while randomly sampling points within the ranges of $a_{1z},a_{2z},\beta$ and $\lambda$. The multiplicative factor on each parameter ensures that the grid is compact enough to avoid unnecessary exploration in low-likelihood regions, yet broad enough to enclose the entire posterior distribution and account for potential systematics.
@@ III Results
We employ the AV sampler (based on the Varaha sampler Tiwari et al. (2023); Tiwari (2024)) to carry out adaptive importance sampling to evaluate marginalized likelihood and to construct posteriors. Since likelihood evaluation is not a bottleneck, we calculate the full Whittle likelihood artist Moran and Whittle (1951) over the frequency range of $0.1$ mHz to $0.1$ Hz at the full Fourier bin width. However, for our analysis of the LDC-1A dataset we integrate up to $0.05$ Hz, which is the Nyquist frequency of that dataset. We confirm the convergence of the algorithm by ensuring that the Jensen-Shannon divergence (JSD) Lin (1991) between the posteriors from the last two iterations is equal to or below $0.001$ bit. JSD quantifies the dissimilarity between two posterior distributions, where a value of 00 bit indicates identical distributions and $1$ bit signifies completely different distributions. A JSD of $0.007$ bit is considered to indicate a significant difference in the posteriors The LIGO Scientific Collaboration et al. (2021a). By ensuring a JSD of less than $0.001$ bit between the last two iterations, we ensure that the interpolation errors are comparable to sampling errors, which are of $O(10^{-4})$ bit Jan et al. (2024). All the posteriors presented here were generated using $20,000$ samples.
@@ III.1 Zero-noise injection
We inject an MBHB signal, with parameters summarized in Table 1, into approximately four months of data sampled at five-second intervals at a zero-noise realization. The signal was generated using numerical relativity surrogate model NRHybSur3dq8, allowing us to incorporate a much larger set of higher-order modes in our analysis compared to previous studies Marsat et al. (2021); Katz et al. (2020), with the injected signal including content from the following modes: (2,2),(3,3),(4,4),(5,5),(2,1),(3,2),(3,1),(4,3) and (4,2). The inference was performed using the same model and with the same mode content. This was done to eliminate the impact of model waveform systematics and isolate the impact of higher modes on data analysis. Both the injected waveform and recovery templates were generated at a $f_{\text{min}}$ of $0.06$ mHz. To ensure higher-order modes are excited and significantly impact SNR, we set the mass ratio of the injection to approximately $1/3$, observed at an $\iota$ of $\pi/4$. The distance was chosen such that the signal will be observed at a high SNR of 1000. For likelihood computations, we use the lisa-data-challenge working group python package to generate the power spectral density (PSD), with the specific noise model used being “SciRDv1”.
@@
Figure 2: Zero-noise injection recovery using NRHybSur3dq8: This figure illustrates the results of our analysis of the four month long zero-noise NRHybSur3dq8 injection. For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. In the top plot, the colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values. For clarity, the reflected sky mode is not included in this plot.
@@
The results of our analysis of this mock signal are presented in Fig 2. Due to the high SNR and the presence of higher mode content, the posteriors for mass and spins are Gaussian-like and peak at the injected values. The typical bias observed in sky localization for (2,2) only mode analyses Marsat et al. (2021) is absent, with the posteriors peaking at the true values. Furthermore, when considering the full time and frequency dependence of the LISA response and conducting a complete inspiral-merger-ringdown analysis, only the true and reflected modes [^1] The reflected sky mode (*) is related to the true sky mode through the following transformations in the LISA ($L$) frame:          $\beta^{*}_{L}=-\beta_{L}$, $\lambda^{*}_{L}=\lambda_{L}$. for sky position are observed. The reflected mode for this signal, located at $\beta=0.436$ and $\lambda=1.199$, has a low probability; for clarity, we exclude the reflected sky mode from Fig 2. Additionally, we note that the degeneracy between $D_{L}$ and $\iota$ Usman et al. (2019) is broken, resulting in two Gaussian-like posteriors that peak at the injected values. In contrast, in Sec. III.2 where the injection and recovery is performed using a (2,2) only mode model, we begin to observe biases in sky location, $D_{L}$, and $\iota$.